OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x^2 = −3x − 10?

Parth (parthkohli):

x - intercept is the value of x when y = 0. Now just set y = 0 and solve it. You'll get a quadratic though.

OpenStudy (chaise):

y − x^2 = −3x − 10? y=x^2-3x-10 In order to determine the x intercept, make y=0 0=x^2−3x − 10? 0=(x-5)(x+2) x=5 x=2 The x intercepts are 5 and 2. I hope this is right ^_^

OpenStudy (anonymous):

Check your intercepts @chaise

OpenStudy (chaise):

5, -2. My bad ^_^. I don't see why you couldn't have corrected me...